### v2

# Policy Contrastive Decoding for Robotic Foundation Models

Abstract
Robotic foundation models, or generalist robot policies, hold immense potential to enable flexible, general-purpose and dexterous robotic systems. Despite their advancements, our empirical experiments reveal that existing robot policies are prone to learning spurious correlations from pre-training trajectories, adversely affecting their generalization capabilities beyond the training data. To tackle this, we propose a novel Policy Contrastive Decoding (PCD) approach, which redirects the robot policy’s focus toward object-relevant visual clues by contrasting action probability distributions derived from original and object-masked visual inputs. As a training-free method, our PCD can be used as a plugin to improve different types of robot policies without needing to finetune or access model weights. We conduct extensive experiments on top of three open-source robot policies, including the autoregressive policy OpenVLA and the diffusion-based policies Octo and $\pi_{0}$. The obtained results in both simulation and real-world environments prove PCD’s flexibility and effectiveness, e.g., PCD enhances the state-of-the-art policy $\pi_{0}$ by 8% in the simulation environment and by 108% in the real-world environment.
1 Introduction
Recent efforts in developing flexible, general-purpose, and dexterous robotic systems have largely focused on robotic foundation models, or generalist robot policies. The overarching goal of such policies is to empower users to instruct robots to perform arbitrary tasks, enabling autonomous task execution with minimal human supervision. Specifically, the policy takes as input a visual observation of the robot’s state combined with a language instruction that defines the task, and outputs a robot action, such as end-effector displacement. Ongoing advancements in scaling real-world robotic data corpora have facilitated the development of numerous robot policies [4, 20, 32, 36], which have demonstrated exceptional effectiveness in controlling various robots across diverse environments and acquiring a wide range of manipulation skills.
Despite their notable achievements, existing robot policies are prone to learning spurious correlations [49, 13, 25] from pre-training trajectories, adversely affecting their ability to generalize outside their training data. As illustrated in Fig. 1, the robot policy predominantly relies on spurious features (e.g., background or textures) rather than object features within observations to predict actions (see (a)(d)). Consequently, slight alterations to the visual observation background, like adjustments to the panning light region (i.e., ($\xrightarrow{}$(b)) and the handle position (i.e., ($\xrightarrow{}$(c)), lead to 36% and 32% drops in the policy’s action prediction performance, respectively. Prior work in other fields has demonstrated that models relying on spurious correlations often suffer significant performance degradation on test data when a distribution shift occurs between the training and test environments [31, 44, 17]. This underscores the importance of redirecting robot policies’ focus from spurious features to object-relevant ones to enhance their generalization across different scenarios during inference. Therefore, we provide a first study on the following question: {mdframed} [skipabove=4pt, innertopmargin=4pt, innerbottommargin=4pt] Can we propose a training-free, plug-and-play approach to mitigate the adverse effects of spurious correlations, thereby rectifying predicted actions for existing robot policies?
In this work, we present a Policy Contrastive Decoding (PCD) approach to answer the question. The core idea of our PCD approach is to shift the robot policy’s attention from spurious features to object-relevant ones by contrasting action probability distributions derived from original and object-masked observation images during inference. To ensure both effectiveness and flexibility of the approach, a Tracking-to-Mask (Track2Mask) strategy is introduced, which first annotates target-specific objects in the initial visual observation using human expertise (e.g., Point and Box prompts) or off-the-shelf object detection models (e.g., Grounding DINO [28]), then employs the SAM2 [21] model to track these objects across subsequent observations in the trajectory. This enables precise object masking with minimal or no human intervention. Moreover, unlike autoregressive policies, diffusion-based policies cannot directly generate action probability distributions to feed into our PCD. To cope with this limitation, we introduce a KDE-based Probabilistic Modeling (KDE-PM) scheme, which calculates action probability distributions for diffusion-based policies using the kernel density estimation (KDE) [47] technique. This enables our PCD approach to be compatible with both autoregressive and diffusion-based policies.
Figure 1: Robot policies tend to spuriously correlate task-irrelevant features with actions, compromising their ability to generalize to unseen scenarios. As observed, changing the light position from (a) to (b) and the drawer handle position from (a) to (c) results in 36% and 32% drops in the performance of the baseline policy OpenVLA [20], respectively. (d) Attention map. More results are in Section 4.4 and Appendix A.1.
To the best of our knowledge, this is the first work to introduce a training-free approach for resolving the spurious correlation issue in robot policies. We conduct extensive evaluations in both simulation and real-world environments, across a total of 15 diverse tasks, and on top of three open-source robot policies—including the autoregressive policy OpenVLA [20] and the diffusion-based policies Octo [32] and $\pi_{0}$ [3]. The achieved results demonstrate the strong flexibility and effectiveness of our PCD approach. On simulation benchmarks, PCD enhances the performance of Octo, OpenVLA and $\pi_{0}$ by 31%, 41% and 8%, respectively. On real-world manipulation tasks, PCD outperforms the state-of-the-art robot policy $\pi_{0}$ by 108%.
Contributions. Our contributions in this work are threefold. 1) We propose PCD, a simple, training-free, and easy-to-implement scheme to tackle the spurious correlation issue in robot policies. 2) PCD can be used as a plugin to enhance both autoregressive and diffusion-based policies, without requiring fine-tuning or access to pre-trained model weights. 3) Extensive experiments demonstrate PCD’s effectiveness and flexibility, consistently improving performance over three state-of-the-art policies across 15 manipulation tasks on multiple robotic platforms.
2 Related Work
Robotic Foundation Models. The rapid development of robotic foundation models or generalist robot policies has been significantly inspired by the success of large visual and language models [18, 52, 42]. Robot policies can be categorized based on action prediction strategies into autoregressive models and diffusion-based policies. Autoregressive policies sequentially generate action tokens based on previous outputs [4, 5, 9, 50, 1, 53, 54]. OpenVLA [20] equips with a Llama 2 [42] language model alongside a visual encoder that merges pretrained features from SigLIP [52] and DINO-v2 [34]. Diffusion-based policies simultaneously generate entire action dimensions by performing diffusion processes on noise [32, 27]. Octo [40] learns a lightweight policy on the OXE dataset [9], enabling rapid adaptation to unknown tasks and scenarios on standard consumer-grade GPUs. $\pi_{0}$ [3] exhibits flexible control over different robot types by leveraging pre-trained general knowledge from vision-language models. Although most robot policies are trained on extensive robotic behavior datasets, recent studies have shown that slight variations in deployment scenarios can significantly affect their generalization performance on downstream tasks in real-world environments [11, 12, 29].
Contrastive Decoding. Recent studies in large vision-language models (VLMs) have highlighted the challenge of hallucination [14, 37, 26], where models generate inaccurate or misleading outputs not grounded in the input data. Contrastive Decoding (CD) [10, 7, 22] suppresses hallucinations by comparing output distributions from original and distorted input without altering the model’s weights. Instruction Contrastive Decoding (ICD) [46] is used to combat hallucinations in multimodal tasks by modulating the confidence in multimodal alignment of the model’s visual and textual inputs, enabling it to distinguish between hallucinated and relevant tokens. Visual Contrastive Decoding (VCD) [35, 10] aims to reduce object hallucinations by comparing outputs from original and distorted visual inputs. This approach provides a computationally efficient solution as it requires neither additional training nor external pre-trained models. Our proposed Policy Contrastive Decoding (PCD) approach in this work is mainly inspired by the success of VCD schemes. However, unlike those schemes that contrast output distributions from original and noise-distorted [35] or image-removed [10] inputs to counteract LVLMs’ overreliance on language priors, our PCD contrasts action probability distributions derived from original and object-masked inputs, thereby redirecting the robot policy’s focus to object-relevant visual features and enhancing their generalization across different scenarios.
3 Proposed Approach
In this section, we present PCD, our training-free, plug-and-play scheme designed to tackle the spurious correlation issue for existing generalist robot policies. Notably, PCD does not require access to the pre-trained parameters of these robot policies; i.e., it treats each robot policy as a black-box.
3.1 Preliminaries
This work focuses on robot manipulation tasks where the objective is to leverage a learned language-conditioned policy $\pi_{\theta}(\boldsymbol{a}_{i}\mid\boldsymbol{o}_{i},\ell)$ to predict an $M$-dimensional action $\boldsymbol{a}_{i}=[a_{1},...,a_{M}]\in\mathbb{R}^{M}$ conditioned on the current visual observation $\boldsymbol{o}_{i}$ and the language instruction $\ell$ of the task. In this part, we formalize two mainstream paradigms for action prediction: autoregressive action prediction (e.g., OpenVLA [20]) and diffusion-based action prediction (e.g., Octo [32] and $\pi_{0}$ [3]).
Figure 2: Overview of our Proposed Policy Contrastive Decoding (PCD) approach. PCD serves as a plugin to redirect the robot policy’s focus toward object-relevant visual cues by contrasting action probability distributions derived from original observations ${p}$ and object-masked observations $\hat{p}$. For illustrative purposes, we visualize the predictions only in the $\Delta x$ and $\Delta y$ dimensions of the robot action space [$\Delta x$, $\Delta y$, $\Delta z$, rotx, roty, rotz, gripper].
Autoregressive Robot Policies. The policy decomposes the action prediction task into sequential token generation. Each action dimension $a_{t}$ of $\boldsymbol{a}_{i}$ is sampled autoregressively from the probability distribution, conditioned on the visual observation $\boldsymbol{o}_{i}$, the language instruction $\ell$, as well as the previously predicted dimensions $a_{<t}=\{{a}_{t}\}_{t=1}^{t-1}$:
$a_{t}\sim\pi_{\theta}\,(a_{t}\mid\ell,\boldsymbol{o}_{i},a_{<t}).$
For example, OpenVLA first encodes $\boldsymbol{o}_{i}$ through a hybrid vision backbone (DINOv2 [6] w/ SigLIP [51] ) and tokenizes $\ell$ using a pre-trained language model (Llama 2 [43]). These multimodal features are fused in the LLM’s embedding space, which then autoregressively emits $M$ discrete tokens. Each token is finally mapped to an action dimension $a_{t}$ via a de-tokenizer [20].
Diffusion-based Robot Policies. The policy frames action generation as an iterative denoising process [15]. Denote $\boldsymbol{e}_{i}$ the multimodal embeddings fusing $\boldsymbol{o}_{i}$ and $\ell$. Starting from Gaussian noise $\boldsymbol{a}_{i}^{K}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, the reverse diffusion process refines all the $M$ dimensions of the action $\boldsymbol{a}_{i}$ in parallel over $K$ steps through:
$\boldsymbol{a}_{i}^{k-1}=\alpha\left(\boldsymbol{a}_{i}^{k}-\gamma\epsilon_{\theta}\left(\boldsymbol{a}_{i}^{k},\boldsymbol{e}_{i},k\right)\right)+\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ (1)
where $\boldsymbol{a}_{i}^{k}$ is the noisy action at diffusion step $k$, $\epsilon_{\theta}\left(\boldsymbol{a}_{i}^{k},\boldsymbol{e}_{i},k\right)$ is the denoising network, and $\alpha,\gamma,\sigma$ are parameters of a cosine noise schedule [33]. By jointly optimizing the entire $M$ action dimensions of $\boldsymbol{a}_{i}$, diffusion-based robot policies enable fine-grained sampling from the joint probability $\pi_{\theta}(\boldsymbol{a_{i}}|l,\boldsymbol{o}_{i})$ of all action dimensions.
3.2 Policy Contrastive Decoding (PCD)
The core idea of our proposed PCD method is to eliminate the adverse effect of spurious correlations by redirecting the robot policy’s focus toward object-relevant features during inference. To facilitate the introduction of our PCD, we define spurious correlations as follows. {mdframed} [skipabove=4pt, innertopmargin=4pt, innerbottommargin=4pt,linecolor=black] Definition 3.1 (Spurious Correlations): Let observations $\boldsymbol{o}$ and language instructions $l$ be generated by a set of $L$ underlying latent factors $\{c_{j}\}_{j=1}^{L}$. These factors are decomposed into task-relevant factors, denoted by the random variable $\boldsymbol{u}$, which exclusively determine the ground-truth expert policy such that $p(\boldsymbol{a}|\boldsymbol{o},l)=p(\boldsymbol{a}|\boldsymbol{u})$, and task-irrelevant factors, denoted by the random variable $\boldsymbol{v}$, which have no causal effect on the expert’s actions $\boldsymbol{a}$. Spurious correlation is then defined as the statistical dependence between the action $\boldsymbol{a}$ and these task-irrelevant factors $\boldsymbol{v}$ observed under the training distribution $p_{\mathrm{train}}(\boldsymbol{a},\boldsymbol{v})$, quantified by the mutual information $I_{\mathrm{train}}(\boldsymbol{a},\boldsymbol{v})$. Spurious correlations, indicated by mutual information $I_{\mathrm{train}}(\boldsymbol{a},\boldsymbol{v})>0$ in the training set, allow a robot policy $\pi_{\theta}$ to infer action $\boldsymbol{a}$ from task-irrelevant factors $\boldsymbol{v}$. These factors act as shortcuts. However, such correlations are unreliable. Under out-of-distribution (OOD) conditions—stemming from environmental shifts or compounding robotic errors—$\boldsymbol{v}$ may change independently of task-relevant factors $\boldsymbol{u}$ and, consequently, $\boldsymbol{a}$, leading to inaccurate action predictions.
A direct approach to mitigate learning from spurious correlations involves filtering out $\boldsymbol{v}$ and utilizing only $\boldsymbol{u}$. Nevertheless, identifying and neutralizing spurious features is challenging due to their significant variability across tasks and scenarios, coupled with interdependencies within the feature space. Drawing inspiration from Vision Contrastive Decoding (VCD) [35, 10], which effectively reduces hallucinations in vision-language models, we introduce Policy Contrastive Decoding (PCD). PCD aims to shift the robot policy’s focus from spurious features $\boldsymbol{v}$ towards task-relevant features $\boldsymbol{u}$ associated with the target object. This is achieved by contrasting model outputs derived from original visual inputs with those from object-masked visual inputs.
Method Overview. Fig. 2 presents an overview of our PCD approach. Suppose we have a pretrained policy $\pi_{\theta}$. Given the current visual observation $\boldsymbol{o}_{i}$ and the language command $\ell$, we first generate two distinct action probability distributions of $\boldsymbol{a}_{i}$: one is $\pi_{\theta}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})$ conditioned on the original visual input $\boldsymbol{o}_{i}$, and the other is $\pi_{\theta}(\boldsymbol{a}_{i}|l,\hat{\boldsymbol{o}}_{i})$ conditioned on object-masked counterpart $\hat{\boldsymbol{o}}_{i}$ that removes all task-relevant factors related to target objects. Then, a new contrastive action probability distribution $\pi_{\theta}^{*}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})$ is obtained by contrasting $\pi_{\theta}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})$ and $\pi_{\theta}(\boldsymbol{a}_{i}|l,\hat{\boldsymbol{o}}_{i})$:
$\pi_{\theta}^{*}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})=\frac{1}{C}\cdot\pi_{\theta}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})\left(\frac{\pi_{\theta}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})}{\pi_{\theta}(\boldsymbol{a}_{i}|l,\hat{\boldsymbol{o}}_{i})}\right)^{\alpha},$ (2)
where $C$ is a normalization constant and a larger value of $\alpha\geq 0$ indicates a stronger amplification of differences between the two distributions. Particularly, $\alpha=0$ recovers the baseline policy. In such a manner, the obtained $\pi_{\theta}^{*}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})$ amplifies model predictions on object-relevant features in $\boldsymbol{u}$, thus exhibiting insensitivity to spurious features in $\boldsymbol{v}$, as illustrated in Fig. 2.
To guarantee the flexibility and practical applicability of our PCD method, we need to address the following two questions: {mdframed} [skipabove=4pt, innertopmargin=4pt, innerbottommargin=4pt] 1) How can object-masked observations $\hat{\boldsymbol{o}}_{i}$ be automatically generated for each trajectory?
2) How can we approximate the policy distribution $\pi_{\theta}$ for diffusion-based policies?
Tracking-to-Mask. We overcome the first question by devising a Tracking-to-Mask (Track2Mask) strategy, which enables precise object masking for sequential visual observations along each trajectory, requiring minimal or no human intervention. Concretely, the first step of Track2Mask is to annotate the task-specified objects in the initial visual observation. Inspired by visual prompting techniques [45, 56], one can annotate the target objects in the initial observation using Point or Box prompts. Besides, we provide the option of leveraging off-the-shelf (open vocabulary) object detection models, such as Grounding DINO [28] for automatic object annotation. Next, the SAM2 [21] model is employed to track and segment the target object across subsequent observations in the trajectory, and the segmented objects are then inpainted to obtain object-masked observations. For more details of Track2Mask, please refer to Appendix A.2.
KDE-based Probabilistic Modeling. We address the second question using KDE-based Probabilistic Modeling (KDE-PM), which approximates action probability distributions for diffusion-based policies through kernel density estimation (KDE) [47]. Specifically, we first use the pretrained diffusion-based policy to sample $N$ candidate action predictions: $\{\boldsymbol{a}_{i}(j)\}_{j=1}^{N}$, where $\boldsymbol{a}_{i}(j)=[{a}_{1}(j),...,{a}_{M}(j)]\sim\pi_{\theta}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})$. With the assumption that all action dimensions are independent, i.e., $\pi_{\theta}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})=\prod\limits_{t=1}^{M}\pi_{\theta}(a_{t}|l,\boldsymbol{o}_{i})$, we calculate an approximated probability distribution for each action dimension of the action $\boldsymbol{a}_{i}$ via KDE:
$\pi_{\theta}(a_{t}|l,\boldsymbol{o}_{i})\approx\frac{1}{C^{\prime}}\sum_{j=1}^{N}\mathcal{K}\left(\frac{a_{t}-a_{t}(j)}{b}\right),\,\,t=1,...,M,$ (3)
where $C^{\prime}$ is a normalization constant and $\mathcal{K}(\cdot)$ indicates a Gaussian kernel $\mathcal{K}(u)=\frac{1}{\sqrt{2\pi}}e^{-\frac{u^{2}}{2}}$, $b$ is the bandwidth parameter controlling the smoothness of the distribution, and $\{a_{t}(j)\}_{j=1}^{N}$ are the $N$ candidate actions for the $t$-th dimension of the action $\boldsymbol{a}_{i}$, generated from the diffusion process. Similarly, we can derive an action probability distribution for $a_{t}$ conditioned on object-masked $\hat{\boldsymbol{o}}_{i}$, i.e., $\pi_{\theta}(a_{t}|l,\hat{\boldsymbol{o}}_{i})$. With the independence assumption, from $\pi_{\theta}(a_{t}|l,\boldsymbol{o}_{i})$ and $\pi_{\theta}(a_{t}|l,\hat{\boldsymbol{o}}_{i})$ we can compute the joint action distributions, which are then used to compute the contrastive action probability $\pi_{\theta}^{*}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})$ using Eq. (2). In this way, our PCD is compatible with diffusion-based robot policies.
Inference with PCD. As a training-free method, our proposed PCD can be employed as a plugin to improve the inference stage of different types of robot policies (i.e., autoregressive policies and diffusion-based policies). Specifically, for a pre-trained robot policy $\pi_{\theta}(\boldsymbol{a}_{i}\mid\boldsymbol{o}_{i},\ell)$, given the current observation ${\boldsymbol{o}}_{i}$, we can obtain the object-masked $\boldsymbol{o}_{i}$ (i.e., $\hat{\boldsymbol{o}}_{i}$) leveraging the devised Track2Mask strategy. Based on $\boldsymbol{o}_{i}$, $\hat{\boldsymbol{o}}_{i}$ and a language command $\ell$, the autoregressive policy directly computes contrastive action probability distributions for deriving $\boldsymbol{a}_{i}$ (w/ Eq. (2)), whereas the diffusion-based policy requires performing the KDE-PM process on its output predictions to achieve the same goal. Pseudocode for our PCD method is provided in Algorithm 1.
Algorithm 1 PCD: Policy Contrastive Decoding
1:A pre-trained robot policy: $\pi_{\theta}(\boldsymbol{a}_{i}\mid\boldsymbol{o}_{i},\ell)$; initial observation $\boldsymbol{o}_{0}$; language command $\ell$; maximum time step $S$.
2:$s\leftarrow 0$
3:while $s\leq S$ do
4:     Obtain object-masked $\boldsymbol{o}_{s}$: $\hat{\boldsymbol{o}}_{s}$triangleright w/ Track2Mask
5:     Obtain $\pi_{\theta}(\boldsymbol{a}_{s}|l,\boldsymbol{o}_{s})$ and $\pi_{\theta}(\boldsymbol{a}_{s}|l,\hat{\boldsymbol{o}}_{s})$triangleright w/ KDE-PM for diffusion-based policies
6:     Compute $\pi_{\theta}^{*}(\boldsymbol{a}_{s}|l,\boldsymbol{o}_{s})$ using Eq. (2)
7:     Sample $\,\boldsymbol{a}_{s}$ from $\pi_{\theta}^{*}(\boldsymbol{a}_{s}|l,\boldsymbol{o}_{s})$ and execute $\boldsymbol{a}_{s}$
8:     $\boldsymbol{o}_{s+1}\leftarrow\text{New observation}$triangleright Obtain a new observation
9:     $s\leftarrow s+1$
10:end while
4 Experiments
In this section, we conduct extensive experiments in both simulation and real-world environments, on top of three state-of-the-art open-source robot policies to answer the following questions: {mdframed} [skipabove=4pt, innertopmargin=4pt, innerbottommargin=4pt] 1) Can PCD improve robot policies in both simulation and real-world environments?
2) How does the performance vary with different design choices?
3) What kinds of spurious correlations can our PCD tackle? We answer the first question in Sections 4.1 and 4.2, the second question in Section 4.3, and the third question in Section 4.4. Detailed descriptions of the baseline robot policies and the evaluation tasks used in simulation and real-world environments are presented in Appendix A.3.
4.1 Simulation Experiments
We perform simulation experiments in SIMPLER [24], a real-to-sim evaluation environment designed specifically for real-robot policies. SIMPLER accurately reflects real-world performance, enabling reliable assessment of robotic policies. We evaluate 9 tasks across two robot platforms: 5 tasks using the Google Robot and 4 tasks using the WidowX arm.
Experimental Setup. We conduct experiments on top of three diverse robot policies, including the autoregressive policy OpenVLA [20], and the diffusion-based policies Octo (base) [32] and $\pi_{0}$ [3]. We treat these policies as black-boxes, applying PCD solely based on the output action probability distributions for OpenVLA and the output actions for Octo and $\pi_{0}$. For our PCD, we annotate the target object in the initial observation using three manners: artificial Point and Box prompts, and automatic detection results from Grounding DINO [28]. The number of the sampled noise vectors $N$ in KDE-PM is set to 24 for Octo and $\pi_{0}$. We perform ablation studies on the hyperparameter $\alpha$ in Eq. (2), as well as on the object detection strategies and object inpainting paradigms of the devised Track2Mask module in Section 4.3.
Results and Analysis. Table 1 reports the performance of three robot policies (i.e., OpenVLA [20], Octo [32] and $\pi_{0}$ [3]) integrated w/ or w/o our PCD across a total of 9 tasks from Google Robot and WidowX. From the results in the table, we have the following key observations. 1) Our PCD consistently improves the three baseline policies over the 9 tasks from both the Google Robot and WidowX robotic platforms. On average, PCD boosts the success rates of the autoregressive policy OpenVLA, and the diffusion-based policies Octo and $\pi_{0}$ by up to 41%, 31% and 8%, respectively. 2) Our PCD exhibits stable and outstanding performance across the three object annotation strategies, including artificial Point and Box prompts and automatic object detection with Grounding DINO, underscoring its effectiveness and adaptability. In particular, on the Octo and OpenVLA baseline policies, leveraging the off-the-shelf model Grounding DINO demonstrates better performance than using human Point and Box prompts, while showing comparable results on $\pi_{0}$, highlighting the practical applicability of the PCD method. 3) Compared to OpenVLA and Octo, the baseline model $\pi_{0}$ demonstrates significantly superior performance across all 9 tasks. For instance, OpenVLA and Octo achieve near-zero success rates on both“Apple Drawer” and “Stack Cube” tasks whereas $\pi_{0}$ achieve success rates of 11% and 67% on these two tasks, respectively. Despite this, from the average results over the 9 tasks, our PCD still enhances the strong baseline $\pi_{0}$ by 8%. In summary, the obtained results in the table demonstrate that our PCD approach can be used as a plugin to improve different types of robot policies.
4.2 Real-world Experiments
In this section, we conduct experiments on real-world tasks to assess PCD’s practical effectiveness.
Table 1: SIMPLER Performance—Success rates of the autoaggressive policy OpenVLA [20] and the diffusion-based policies Octo [32] and $\pi_{0}$ [3] integrated w/ or w/o PCD on 9 tasks from Google Robot and WidowX. Task-specific objects in the initial observation are annotated by artificial Point and Box prompts or the automatic detection results of Grounding DINO [28]. As a plug-and-play approach, PCD consistently enhances the three policies by large margins over the 9 tasks.
OpenVLA[20] Octo [32] $\pi_{0}$ [3]
Tasks Base +PCD Base +PCD Base +PCD
Point Box GDINO Point Box GDINO Point Box GDINO
Average 0.17 0.24 0.21 0.24 0.16 0.21 0.20 0.19 0.65 0.70 0.69 0.69
+41% +24% +41% +31% +25% +19% +8% +6% +6%
Google Robot Close Drawer 0.41 0.68 0.66 0.77 0.33 0.28 0.38 0.16 0.75 0.72 0.76 0.71
Move Near 0.58 0.62 0.52 0.53 0.07 0.13 0.12 0.13 0.73 0.74 0.71 0.75
Open Drawer 0.20 0.29 0.29 0.26 0.00 0.01 0.00 0.02 0.39 0.49 0.50 0.52
Pick Coke Can 0.25 0.40 0.42 0.42 0.28 0.54 0.49 0.52 0.84 0.88 0.87 0.87
Apple Drawer 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.11 0.29 0.24 0.22
WidX. Carrot Plate 0.00 0.08 0.00 0.04 0.14 0.28 0.22 0.19 0.59 0.67 0.62 0.62
Eggplant Basket 0.01 0.02 0.02 0.11 0.48 0.41 0.38 0.44 0.93 0.86 0.87 0.88
Spoon Towel 0.00 0.02 0.00 0.03 0.10 0.15 0.14 0.15 0.83 0.89 0.88 0.86
Stack Cube 0.04 0.01 0.00 0.00 0.00 0.05 0.06 0.07 0.67 0.77 0.78 0.75
Figure 3: Real-world Performance—Success rates and time costs of the state-of-the-art robot policy $\pi_{0}$ [3] integrated w/ or w/o PCD on 6 real-world tasks. The target objects in the initial observation are automatically annotated by Grounding DINO [28]. PCD delivers a remarkable 108% performance improvement on the baseline policy, though it incurs a 24% increase in time cost.
Experimental Setup. The simulation experiments show that $\pi_{0}$ significantly outperforms Octo and OpenVLA. Therefore, we employ $\pi_{0}$ as the baseline policy to validate the effectiveness of our PCD approach, as shown in Fig. 3. Considering the generalization limitations of existing robot policies in real-world environments, we first finetune $\pi_{0}$ on downstream real-world tasks before evaluating PCD’s performance. We design 6 manipulation tasks that focus on evaluating the policy’s performance along multiple dimensions: spatial reasoning and interacting with various objects and scenes. The $\pi_{0}$ policy is jointly fine-tuned on the 6 real-world tasks, with each task consisting of 10 training demonstrations. During inference, we conduct 20 trials for each real-world task and randomize the configurations and orientations of task-specific objects for each trial. We use an AGILEX PIPER 6DOF robot arm in our experiments. For our PCD, we leverage the off-the-shelf model Grounding DINO to detect task-specific objects in the initial observation, and use LaMa [39] to inpaint those objects segmented by the SAM2 [21] model. All other hyperparameters are kept consistent with the settings applied in the simulation experiments.
Results and Analysis. Fig. 3 reports the success rates as well as time costs of the state-of-the-art robot policy $\pi_{0}$ [3] integrated w/ or w/o our PCD on 6 real-world tasks. For each task, we record the time cost (in seconds) from the start of the inference process to the successful completion of the task. From the obtained results in the figure, we have several key observations. 1) PCD consistently improves the success rates of the strong baseline policy across six real-world tasks, achieving an average enhancement of 108%. 2) While PCD significantly boosts success rates, it also increases the time cost by 24% on average compared to the baseline policy. However, considering the notable improvement in task completion success rates, this trade-off is acceptable in a wide range of real-world robotic applications. 3) PCD demonstrates consistent performance improvements across tasks of varying complexity. For example, in the challenging "Stack Cube" task, PCD improves the success rate from 0.05 to 0.10, showing its strong adaptability and robustness across diverse scenarios. Additionally, the backgrounds of these six real-world tasks are far more complex than those of the nine simulation tasks, containing numerous distracting elements such as plastic bottles, baskets, and trash cans. Nevertheless, our PCD shows remarkable advantages. It is worth mentioning that we use consistent PCD hyperparameters across tasks in both the simulation and real-world environments, without carefully tuning task-specific PCD hyperparameters (e.g., $\alpha$) for each individual task, leaving room for potential performance improvements through task-specific parameter tuning.
4.3 Ablation Studies
In this section, we conduct ablation studies to explore how the performance of our PCD varies with different design decisions. We perform experiments in the SIMPLER simulation environment, and report the average results across the nine tasks for our PCD applied to the three baseline policies.
Effect of $\alpha$. As formulated in Eq. (2), PCD leverages the hyperparameter $\alpha$ to control the level of amplification between output distributions from original and object-masked visual inputs. A larger $\alpha$ value indicates a stronger amplification of differences between the two distributions, while $\alpha=0$ reduces to regular prediction. We adjust $\alpha$ by setting it to {0, 0.2, 0.4, 0.6, 0.8, 1.0}, and report the performance of our PCD method over the 9 tasks in Fig. 4 (a). $\alpha$ = 0 represents the baseline policies. As can be observed, our PCD consistently improves Octo, OpenVLA and $\pi_{0}$ when $\alpha>0$, demonstrating its effectiveness and stability. Particularly, PCD yields the best results on Octo, OpenVLA and $\pi_{0}$ when the values of $\alpha$ are set to 1.0, 0.8, and 0.2, respectively.
Figure 4: Ablation studies on (a) the hyperparameter $\alpha$ in Eq. (2), (b) the object detection schemes and (c) object inpainting strategies in Track2Mask. $\alpha=0$ in (a) and the black dotted lines in (b)(c) represent the performance of the baseline policies. The results are averaged over the 9 simulation tasks. PCD consistently improves the three policies when $\alpha>0$ and exhibits low sensitivity to changes in off-the-shelf object detection and inpainting strategies.
Effect of Off-the-shelf Object Detection Models. Beyond leveraging artificial Point and Box prompts, the devised Track2Mask module can employ off-the-shelf open vocabulary object detection models to annotate task-specific objects in the initial observation of the trajectory during inference. Fig. 4 (b) presents an ablation study on various automatic object detection models, including the Grounding DINO [28] model discussed in the main paper, as well as the open vocabulary object detection model YOLO World [8] and semantic segmentation model SED [48]. The output of these models is fed into the SAM2 [21] model for tracking and segmenting the detected objects across subsequent observations in the trajectory. As can be seen, our PCD consistently improves the baseline policies when integrated with each of the three off-the-shelf models, showcasing its effectiveness and stability. Generally, the Grounding DINO model yields the best results among the three models.
Effect of Object Inpainting Strategies. Once the objects specified in the language instruction of the target task are segmented by the SAM2 model, they are inpainted to create object-masked observations. In this experiment, we investigate the effect of different object inpainting strategies on PCD’s performance. Fig. 4 (c) shows the results of PCD integrated with three inpainting strategies: Telea [41] and Navier-Stokes [2] and LaMa [39]. As indicated in the figure, our PCD approach consistently improves the baseline policies across all strategies, showcasing its effectiveness and stability. Furthermore, PCD demonstrates low sensitivity to variations in the three object inpainting strategies, further validating its applicability and robustness. Notably, LaMa delivers the best results across all three baseline policies.
4.4 Robustness to Various Kinds of Spurious Correlations
The central concept of our proposed PCD approach in this work is to overcome the adverse effects of spurious features on the decision-making of robot policies. So, what types of spurious correlations can PCD effectively address? To answer the question, we evaluate the robustness of the OpenVLA [20] and $\pi_{0}$ [3] policies to task scene variations in both simulation and real-world environments. As illustrated in Fig. 5, we report the performance of the two baseline policies integrated w/ or w/o our proposed PCD approach in different testing scenarios with unseen spatial relationships, brightness, distractors and table textures within visual observations. The results shown in the figure lead to the following observations. 1) Consistent with the previous experimental results, our PCD approach significantly improves the performance of the baseline policies on both simulation and real-world tasks (Orange line vs. Blue line). 2) Both baseline policies experience substantial drops in performance when testing scenarios are modified. For example, altering the brightness of visual observations results in 48% and 75% decreases in success rates for the two policies, respectively. 3) PCD consistently alleviates the performance degradation of the baseline policies across all unseen testing scenarios. Remarkably, after integrating PCD, the two baseline policies even performs better in 4/10 of unseen scenarios than in the original training scenarios. The results in the figure reveal that our proposed PCD approach can serve as a plugin to mitigate the adverse effects of various types of spurious correlations on the generalization of robot policies.
[…]
5 Conclusion, Limitations and Future Work
In this work, propose Policy Contrastive Decoding (PCD) to tackle the adverse effects of spurious correlations learned by generalist robot policies. PCD redirects the policy’s attention from spurious features to object-relevant ones during inference by contrasting action probability distributions obtained from original and object-masked inputs. As a plug-and-play approach, PCD can enhance both autoregressive and diffusion-based policies, without requiring fine-tuning or access to pre-trained model weights. Comprehensive experiments demonstrate PCD’s flexibility and effectiveness, achieving consistent improvements over three state-of-the-art robot policies across a total of 15 tasks in simulation and real-world environments.
While our PCD demonstrates effectiveness and flexibility, there are certain limitations. 1) PCD performs object detection and segmentation for each visual observation in the trajectory to produce object-masked observations, which increases the computational overhead of the baseline police. Recent progress in fast LLM inference techniques [55, 23, 30] could potentially enhance the computational efficiency of our PCD approach. 2) This work concentrates exclusively on addressing the spurious correlation issue in robot policies during the testing stage, without exploring methods to prevent the learning of such correlations during training. Hence, future work will focus on investigating effective strategies to tackle spurious correlations at both the training and inference levels.
Appendix A Appendix
A.1 CAM Visualization of Failure Cases.
We present the CAM [38] visualization of failure cases for the OpenVLA [^1] We only visualize the CAM results of OpenVLA, as the action prediction mechanism of diffusion-based models (e.g., Octo and $\pi_{0}$) makes it difficult to produce their CAM results. model across the nine tasks in Fig. 6. For each task, the averaged CAM results over the seven action dimensions ([$\Delta x$, $\Delta y$, $\Delta z$, rotx, roty, rotz, gripper]) are plotted. As can be observed from the figure, the model tends to predict actions based on spurious or background features in visual observations. For instance, on the “move the blue plastic bottle near sponge” (denoted as “Move Near”) task, the model model habitually focuses on the “coke can” object that it frequently encountered during the training phase. These observations demonstrate that the learned policy utilizes perceptual data for decision-making in a way that significantly diverges from human strategies. Robot policies often erroneously associate task-irrelevant features with actions, a major cause of overfitting and poor generalization across tasks.
Figure 6: CAM visualization of failure cases.
A.2 Details of the Track2Mask Module
In our devised Track2Mask module, the target object specified by the language instruction in the initial observation can be annotated using Point and Box prompts, along with off-the-shelf open-vocabulary object detection models, as illustrated in Fig. 7. The SAM2 [21] model is then applied to automatically track and segment the target object across subsequent observations in the trajectory, which enables precise object masking with minimal or no human intervention during inference. Note that, once the task-specific objects are segmented by the SAM2 model, they are inpainted to create object-masked observations.
Figure 7: Illustration of the Track2Mask module.
A.3 Details of the Experimental Setup
A.3.1 Baseline Policies
Simulation Experiments. We conduct simulation experiments using three diverse robot policies in the SIMPLER environment, including the autoregressive policy OpenVLA [20], and diffusion-based policies Octo (base) [32] and $\pi_{0}$ [3].
  • OpenVLA: a vision-language-action model with 7 billion parameters, is trained on 970,000 episodes of robotic demonstrations from the Open X-Embodiment dataset. This policy is fine-tuned using the pre-trained Prismatic [19] model. [^2] https://openvla.github.io/
  • Octo (base): an open-source generalist policy with 93 million parameters, is pre-trained on a blend of 25 different datasets from the Open X-Embodiment dataset. It utilizes a transformer backbone derived from ViT-B, accompanied by a diffusion action head to model expressive action distributions. [^3] https://octo-models.github.io/
  • $\pi_{0}$: a flow matching architecture built on top of a pre-trained vision-language model to leverage Internet-scale semantic knowledge. To facilitate evaluation in the SIMPLER environment, we leverage the reproduced version of $\pi_{0}$, pretrained on tasks from this environment. [^4] https://github.com/allenzren/open-pi-zero
Real-world Experiments. Since the simulation experiments show that $\pi_{0}$ significantly outperforms Octo and OpenVLA, we conduct real-world experiments using the $\pi_{0}$ policy  [^5] https://www.physicalintelligence.company/blog/pi0. The policy is finetuned using on six real-world tasks, with each task consisting of 10 trajectories. We employ the parameter-efficient finetuning strategy LoRA [16] to adapt $\pi_{0}$ to these real world tasks. The training hyperparameters are listed in Table 2.
Table 2: Hyperparametes for finetuning $\pi_{0}$ on real-world tasks.
Hyperparametes Setting
Batch Size 128
Optimizer AdamW
Learning Rate 2.5e-5
LR Schedule Cosine Decay
Weight Decay 1e-10
Training Step 25000
Action Chunk 10
LoRA Rank 16
A.3.2 Evaluation Tasks
We conduct extensive evaluations in both simulation and real-world environments using a total of 15 diverse tasks. The details of these tasks are presented in Table 3. We conduct 20 trials for each real-world task, randomizing the task configurations in every trial, as illustrated in Fig. 8.
Table 3: Details of simulation and real-world tasks.
Env. Tasks Language Instruction Task-specific Objects # Trials
Simulation Close Drawer close top/middle/bottom drawer top/middle/bottom drawer 100
Move Near move X near Y apple, orange, pepsi can, … 100
Open Drawer open top/middle/bottom drawer top/middle/bottom drawer 100
Pick Coke Can pick coke can coke can 100
Apple Drawer open top drawer/place apple into top drawer top drawer, apple 100
Carrot Plate put carrot on plate carrot, plate 100
Eggplant Basket put eggplant into yellow basket eggplant, yellow basket 100
Spoon Towel put the spoon on the towel spoon, towel 100
Stack Cube stack the green block on the yellow block green cube, yellow cube 100
Real-world Pick Ball pick the ball ball 20
Pick Plug pick the plug plug 20
Move Near move the can near the apple/watermelon can, apple, watermelon 20
Cookies Towel put the cookies on the towel cookies, towel 20
Banana Plate put the banana in the plate banana, plate 20
Stack Cube stack the green cube on the yellow cube green cube, yellow cube 20
Figure 8: We conduct 20 trials for each real-world task, randomizing the task configurations in every trial. The language command for this task is “stack the green cube on the yellow cube.”
A.4 Performance at the Maximum Step
In Table 1, tasks completed before the predefined maximum step are included in the success rate calculation, and the task is terminated upon completion. Table 4 presents success rates computed at the maximum step, irrespective of whether tasks were completed earlier. As seen, PCD demonstrates its advantages by improving the success rates of OpenVLA, Octo and $\pi_{0}$ by up to 33%, 14% and 7%, respectively. We also observe that the vast majority of models underperform on these tasks relative to their results in Table 1, which indicates that even though they complete the task successfully ahead of the maximum steps, their later predicted actions lead to task failure. The superiority of the proposed PCD method is compellingly evidenced by the consistent performance gains over baseline policies presented in Table 1 and Table 4.
Table 4: Success rates (%) of three baseline policies integrated w/ or w/o our PCD over 9 tasks from Google Robot and WidowX—calculated at the maximum step.
OpenVLA[20] Octo [32] $\pi_{0}$ [3]
Model Base +PCD Base +PCD Base +PCD
Point Box DINO Point Box DINO Point Box DINO
Average 0.15 0.20 0.20 0.20 0.14 0.16 0.15 0.16 0.60 0.64 0.64 0.63
+33% +33% +33% +14% +7% +14% +7% +7% +5%
Google Robot Close Drawer 0.41 0.65 0.66 0.76 0.31 0.27 0.37 0.16 0.75 0.71 0.76 0.70
Move Near 0.46 0.46 0.43 0.33 0.05 0.10 0.09 0.12 0.66 0.64 0.62 0.66
Open Drawer 0.20 0.27 0.27 0.23 0.00 0.01 0.00 0.02 0.34 0.44 0.47 0.48
Pick Coke Can 0.24 0.37 0.41 0.37 0.23 0.47 0.38 0.46 0.80 0.86 0.83 0.84
Apple Drawer 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.11 0.29 0.24 0.22
WidX. Carrot Plate 0.00 0.02 0.00 0.03 0.10 0.18 0.16 0.12 0.56 0.62 0.57 0.56
Eggplant Basket 0.01 0.01 0.02 0.08 0.47 0.32 0.28 0.40 0.91 0.80 0.80 0.83
Spoon Towel 0.00 0.01 0.00 0.01 0.09 0.08 0.09 0.08 0.81 0.87 0.86 0.83
Stack Cube 0.00 0.01 0.00 0.00 0.00 0.03 0.02 0.05 0.48 0.57 0.59 0.57